# BoQ: A Place is Worth a Bag of Learnable Queries

### v1

@@ 1 Introduction
Visual Place Recognition (VPR) consists of determining the geographical location of a place depicted in a given image, by comparing its visual features to a database of previously visited places. The dynamic and ever-changing nature of real-world environments pose significant challenges for VPR [34, 58]. Factors such as varying lighting conditions, seasonal changes and the presence of dynamic elements such as vehicles and pedestrians introduce considerable variability into the appearance of a place. Additionally, changes in viewpoint and image scale can expose previously obscured areas, further complicating the recognition process. These challenges are exacerbated by the operational constraints of VPR systems, which often need to operate in real-time and under limited memory. Consequently, there is a compelling need for efficient algorithms capable of generating compact yet robust representations.
@@
With the rise of deep learning, numerous VPR-specific neural networks have been proposed [6, 29, 42, 31, 21, 3, 2, 11, 4], leveraging Convolutional Neural Networks (CNNs) to extract high-level features, followed by aggregation layers that consolidate these features into a single global descriptor. Such end-to-end trainable architectures have been instrumental in enhancing the efficiency and performance of VPR systems.
@@
Recently, Vision Transformers (ViT) [19] have demonstrated remarkable performance in various computer vision tasks, including image classification [15], object detection [14, 32] and semantic segmentation [59]. Their success can be attributed to their self-attention mechanism, which captures global dependencies between distant parts of the image [22]. However, current Transformer-based VPR techniques [47, 60, 55, 50], often rely on reranking [8], a post-processing step aimed at refining the initial set of candidate locations identified through global descriptor search. The reranking process is usually done with geometric verification (e.g. RANSAC) on stored local patch tokens, which is computationally and memory intensive. Moreover, the global retrieval process in Transformer-based methods, whether through specific aggregation of local patches [55, 47] or using the class token of the ViT [60], has yet to reach the performance levels of non-Transformer-based approaches like MixVPR [4] and CosPlace [59].
@@
In this paper, we bridge this performance gap, by introducing a novel Transformer-based aggregation technique, called Bag-of-Queries (BoQ), that learns a set of embeddings (global queries) and employs a cross-attention mechanism to probe local features coming from the backbone network. This approach enables each global query to consistently seek relevant information uniformly across input images. This is in contrast with self-attention-based techniques [47, 55], where the queries are dynamically derived from the input itself. Furthermore, BoQ is designed for end-to-end training, thus seamlessly integrating with both conventional CNN and ViT backbones. Its effectiveness is validated through extensive experiments on multiple large-scale benchmarks, consistently outperforming state-of-the-art techniques, including MixVPR [4], EigenPlaces [12], and NetVLAD [6]. More importantly, BoQ, a single-stage (global) retrieval method that does not employ reranking, outperforms two-stage retrieval methods like TransVPR [47] and R2Former [60], while being orders of magnitude more efficient in terms of computational and memory resources.
@@ 2 Related Works
Early visual place recognition (VPR) techniques relied on hand-crafted local features such as SIFT [33], SURF [9] and ORB [41], which were aggregated into global descriptors using techniques like Bag-of-Words (BoW) [39, 45, 20] or Vector of Locally Aggregated Descriptors (VLAD) [27, 5]. BoW involves learning a visual vocabulary (or set of clusters), where each visual word (or cluster) represents a specific visual characteristic. Images are then represented as histograms, with bins indicating the frequency of each visual word. VLAD extends on this by capturing first-order statistics by accumulating the differences between local descriptors and their respective cluster centers.
@@
CNN-based VPR. The advent of deep learning marked a significant shift in VPR techniques, with various aggregation architectures proposed. Arandjelovic et al. [6] introduced NetVLAD, a trainable version of VLAD that integrates with CNN backbones, achieving superior performance over traditional methods. Following its success, many researchers proposed variants such as CRN [29], SPE-VLAD [53], Gated NetVLAD [56] and SSL-VLAD [36]. Other approaches focus on regions of interests within feature maps [7, 44, 57]. Another key aggregation method in image retrieval is the Generalized Mean (GeM) [40], which is a learnable form of global pooling. Building on GeM, Berton et al. [59] have recently introduced CosPlace, enhancing GeM by integrating it with a linear projection layer, achieving remarkable performance for VPR. Ali-bey et al. [4] introduced MixVPR, an all-MLP aggregation technique that achieved state-of-the-art performance on multiple benchmarks. Another facet of VPR is the training procedure, where the focus is on the loss function [3, 2, 12, 30].
@@
Transformer-based VPR. The Transformer architecture was initially introduced for natural language processing [46], and later adapted to Vision Transformers (ViT) for computer vision applications [19]. In VPR, they have recently been used as backbone combined with MixVPR [26, 25] or NetVLAD [49] resulting in a performance boost compared to CNN backbones. Furthermore, AnyLoc [28] used features extracted from a foundation model, called DINOv2 [37], combined with unsupervised aggregation methods such as VLAD, resulting in notable performance on a wide range of benchmarks.
@@
Recently, two-stage retrieval strategy has gained prominence in Transformers-based VPR. It starts with global retrieval using global descriptors to identify top candidates, followed by a computationally-intensive reranking phase that refines the results based on local features. In this context, Wang et al. [47] introduced TransVPR, which uses CNNs for feature extraction and Transformers for attention fusion to create global image descriptors, with additional patch-level descriptors for geometric verification. On the other hand, Zhang et al. [55] proposed ETR, a Transformer-based reranking technique that uses a CNN backbone for local and global descriptors. Their method leverages cross-attention between the local features of two images to cast the reranking as classification. Recently, Zhu et al. [60] proposed a unified framework integrating global retrieval and reranking, solely using transformers. For global retrieval, it employs the class token and utilizes other image tokens as local features for the reranking module. These two-stage techniques showed great performance, but at the expense of more computation and memory overheads. However, their global retrieval performance is still very limited compared to one-stage methods such as MixVPR [4] and CosPlace [11]. In this work, we propose, Bag-of-Queries (BoQ), a Transformer-based aggregation technique for global retrieval, that outperforms existing state-of-the-art without relying on reranking. This makes BoQ particularly suitable for applications where computational resources are limited, yet high accuracy and efficiency are essential.
@@ 3 Methodology
In visual place recognition, effective aggregation of local features is crucial for accurate and fast global retrieval. To address this, we propose the Bag-of-Queries (BoQ) technique, a novel aggregation architecture that is end-to-end trainable and surprisingly simple, as depicted in Fig. 2.
Our technique rely on Multi-Head Attention (MHA) mechanism [46], which takes three inputs, queries (), keys () and values (), and linearly project them into multiple parallel heads. The output, for each head, is computed as follows:
@@
This ensures that the final descriptor combines information aggregated at different levels of the network, forming a rich and hierarchical representation. The final step involves reducing the dimensionality of the output using one or two successive linear projections with weights and , similar to the approach in [3, 4].
This is followed by -normalization, to bring the global descriptor to the unit hypersphere, which optimizes the similarity search. The training is performed using pair-based (or triplet-based) loss functions that are widely used in the literature [3].
Relations to other methods. The Bag-of-Queries (BoQ) approach bears conceptual resemblance to Detection Transformer (DETR) model [14] in its use of object queries. However, BoQ is fundamentally different, as its global queries are exclusively used to extract and aggregate local feature information from the input; these queries do not contribute directly to the final representation, as demonstrated by the absence of any residual connection between the global queries and the output of the cross-attention. This is in contrast to DETR, where the object queries are integral part to the output, upon which object detection and classification is done.
In comparison to NetVLAD [6], which employs a set of learned cluster centers, and aggregates local features by calculating the weighted sum of their distances to each center. BoQ, leverages cross-attention between its learned queries and the input local feature to dynamically assess their relevance.
@@ 4.1 Datasets
Our experiments are performed on existing benchmarks that present a wide range of real-world challenges for VPR systems. Tab. 1 provides a brief summary of these benchmarks. MapillarySLS (MSLS) [51] was collected using dashcams and features a broad range of viewpoints and lighting changes, testing the system’s adaptability to typical variations in ground-level VPR. Pitts250k [45] and Pitts30k datasets, extracted from Google Street View, exhibit significant changes in viewpoint, which tests the system’s ability to maintain recognition across varying angles and perspectives. AmsterTime [52] presents a unique challenge with historical grayscale images as queries and contemporary color images as references, covering temporal variations over several decades. Eynsham [17] consists entirely of grayscale images, adding complexity due to the absence of color information. Nordland [43] dataset was collected over four seasons, using a camera mounted on the front of a train. It encompasses scenes ranging from snowy winters to sunny summers, presenting the challenge of extreme appearance changes due to seasonal variations. Note that two variants of Nordland are used in VPR benchmarking: one uses a subset of 2760 summer images as queries against the entire winter sequence as reference images (marked with *), while the other uses the entire winter sequence as query images against the entire summer sequence as reference images (marked with **). SVOX [10] stands out with its comprehensive coverage of weather conditions, including overcast, rainy, snowy, and sunny images, to test adaptability to meteorological changes. Additionally, SVOX Night subset focuses on nocturnal scenes, a challenging scenario for most VPR systems.
@@
| MSLS [51] | k | ||||
| Pitts250k [45] | k | k | |||
| Pitts30k [45] | k | k | |||
| AmsterTime [52] | |||||
| Eynsham [17] | k | k | |||
| Nordland* [54] | k | ||||
| Nordland** [43] | k | k | |||
| St-Lucia [35] | |||||
| SVOX [10] | k | k | |||
| SPED [54] | |||||
@@ 4.2 Implementation details
Architecture. For our experiments, we primarily employ pre-trained ResNet-50 backbones for feature extraction [24]. Importantly, our proposed Bag-of-Queries (BoQ) technique is architecturally agnostic and seamlessly integrates with various CNNs or Vision Transformer (ViT) backbones. When employing a ResNet, we crop it at the second last residual to preserve a higher resolution of local features. Additionally, we freeze all but the last residual block and add a convolution to reduce the number of channels, which decreases the computational and memory footprint of the self-attention mechanism. Implementing the BoQ model is straightforward with popular deep learning frameworks such as PyTorch [38] and TensorFlow [1], both providing highly optimized implementations of self-attention and cross-attention, which are the basic blocks of our technique.
Training. We train BoQ following the standard framework of GSV-Cities [3], which provides a highly accurate dataset of k images depicting k places. For the loss function, we use Multi-Similarity loss [48], as it has been shown to perform best for visual place recognition [3]. We use batches containing places, each depicted by images resulting in mini-batches of images. We use the AdamW optimizer with a learning rate of depending the batch-size, and a weight decay of . The learning rate is multiplied by after each epochs. Finally, we train for a maximum of epochs (including epochs of linear warmup) using images resized to . Note that while current techniques such as NetVLAD [6], CosPlace [59] and EigenPlaces [12] train on images of size —effectively triple the pixel count of our chosen resolution—our BoQ model still achieves better performance, underscoring its effectiveness.
Evaluation metrics. We follow the same evaluation metric of existing literature [6, 29, 51, 54, 47, 11, 4, 3, 2, 12], where the recall@ is measured. Recall@ is defined as the percentage of query images for which at least one of the top- predicted reference images falls within a predefined threshold distance, which is typically meters. However, for the Nordland benchmark, which features aligned sequences, the threshold is frame counts for Nordland** and only frame for Nordland*, which makes the latter more challenging.
@@ 4.3 Comparison with previous works
In this section, we present an extensive comparison of our proposed method, Bag-of-Queries (BoQ), with a wide range of existing state-of-the-art VPR aggregation techniques. This includes global average pooling (AVG) [6], Generalized Mean (GeM) [40], NetVLAD [6] alongside its recent variants, SPE-NetVLAD [53] and Gated NetVLAD [56]. Importantly, we compare against recent cutting-edge techniques, such as ConvAP [3], CosPlace [59], EigenPlaces [12] and MixVPR [4], which currently hold the best scores in most existing benchmarks, setting a high standard for BoQ to measure up against in the field of VPR.
Moreover, despite BoQ being a global retrieval technique which does not employ reranking, we compare it with current state-of-the-art two-stage retrieval techniques [18, 13, 23, 47, 60]. These methods leverage geometric verification to enhance recall@k performance at the expense of increased computation and memory overhead.
Results analysis. In Tab. 2, we present a comparative analysis across several challenging datasets, focusing on Recall@1 (R@1), Recall@5 (R@5), and Recall@10 (R@10). We also offer insight into the performance of each method with regard to the top-ranked retrieved images.
@@
In Tab. 3, we follow the same evaluation style proposed by Berton et al. [12], where benchmarks are categorized into multi-view datasets, where images are oriented in various angles, and frontal-view dataset where images are mostly forward facing. Top-1 retrieval performance (R@1) is reported for each technique.
@@
On SVOX, BoQ achieves best results under all conditions, especially on SVOX night, where it achieves R@1, outperforming the second best method (MixVPR) by percentage points. This highlights BoQ’s potential under low-light conditions.
Furthermore, we compare BoQ to two-stage retrieval techniques in Tab. 4, by reporting performance on MSLS-val and Pitts30k-test. Although our technique does not perform reranking, its global retrieval performance surpasses that of existing two-stage techniques, including Patch-NetVLAD (2021), TransVPR (2022) and R2Former (2023). Importantly, this makes our approach more efficient in terms of memory and computation: over faster retrieval time compared to R2Former, which is currently the fastest two-stage VPR technique.

### v2

# : A Place is Worth a Bag of Learnable Queries

@@ 1 Introduction
Visual Place Recognition (VPR) consists of determining the geographical location of a place depicted in a given image, by comparing its visual features to a database of previously visited places. The dynamic and ever-changing nature of real-world environments pose significant challenges for VPRĀ [35, 60]. Factors such as varying lighting conditions, seasonal changes and the presence of dynamic elements such as vehicles and pedestrians introduce considerable variability into the appearance of a place. Additionally, changes in viewpoint and image scale can expose previously obscured areas, further complicating the recognition process. These challenges are exacerbated by the operational constraints of VPR systems, which often need to operate in real-time and under limited memory. Consequently, there is a compelling need for efficient algorithms capable of generating compact yet robust representations.
@@
With the rise of deep learning, numerous VPR-specific neural networks have been proposedĀ [6, 30, 43, 32, 22, 3, 2, 11, 4], leveraging Convolutional Neural Networks (CNNs) to extract high-level features, followed by aggregation layers that consolidate these features into a single global descriptor. Such end-to-end trainable architectures have been instrumental in enhancing the efficiency and performance of VPR systems.
@@
Recently, Vision Transformers (ViT)Ā [20] have demonstrated remarkable performance in various computer vision tasks, including image classificationĀ [15], object detectionĀ [14, 33] and semantic segmentationĀ [61]. Their success can be attributed to their self-attention mechanism, which captures global dependencies between distant parts of the imageĀ [23]. However, current Transformer-based VPR techniquesĀ [49, 62, 57, 52], often rely on rerankingĀ [8], a post-processing step aimed at refining the initial set of candidate locations identified through global descriptor search. The reranking process is usually done with geometric verification (e.g. RANSAC) on stored local patch tokens, which is computationally and memory intensive. Moreover, the global retrieval process in Transformer-based methods, whether through specific aggregation of local patchesĀ [57, 49] or using the class token of the ViTĀ [62], has yet to reach the performance levels of non-Transformer-based approaches like MixVPRĀ [4] and CosPlaceĀ [61].
@@
In this paper, we bridge this performance gap, by introducing a novel Transformer-based aggregation technique, called Bag-of-Queries (BoQ), that learns a set of embeddings (global queries) and employs a cross-attention mechanism to probe local features coming from the backbone network. This approach enables each global query to consistently seek relevant information uniformly across input images. This is in contrast with self-attention-based techniquesĀ [49, 57], where the queries are dynamically derived from the input itself. Furthermore, BoQ is designed for end-to-end training, thus seamlessly integrating with both conventional CNN and ViT backbones. Its effectiveness is validated through extensive experiments on multiple large-scale benchmarks, consistently outperforming state-of-the-art techniques, including MixVPRĀ [4], EigenPlacesĀ [12], and NetVLADĀ [6]. More importantly, BoQ, a single-stage (global) retrieval method that does not employ reranking, outperforms two-stage retrieval methods like TransVPRĀ [49] and R2FormerĀ [62], while being orders of magnitude more efficient in terms of computational and memory resources.
@@ 2 Related Works
Early visual place recognition (VPR) techniques relied on hand-crafted local features such as SIFTĀ [34], SURFĀ [9] and ORBĀ [42], which were aggregated into global descriptors using techniques like Bag-of-Words (BoW)Ā [40, 46, 21] or Vector of Locally Aggregated Descriptors (VLAD)Ā [28, 5]. BoW involves learning a visual vocabulary (or set of clusters), where each visual word (or cluster) represents a specific visual characteristic. Images are then represented as histograms, with bins indicating the frequency of each visual word. VLAD extends on this by capturing first-order statistics by accumulating the differences between local descriptors and their respective cluster centers.
@@
CNN-based VPR. The advent of deep learning marked a significant shift in VPR techniques, with various aggregation architectures proposed. ArandjelovicĀ et al.Ā [6] introduced NetVLAD, a trainable version of VLAD that integrates with CNN backbones, achieving superior performance over traditional methods. Following its success, many researchers proposed variants such as CRNĀ [30], SPE-VLADĀ [55], GatedĀ NetVLADĀ [58] and SSL-VLADĀ [37]. Other approaches focus on regions of interests within feature mapsĀ [7, 45, 59]. Another key aggregation method in image retrieval is the Generalized Mean (GeM)Ā [41], which is a learnable form of global pooling. Building on GeM, BertonĀ et al.Ā [61] have recently introduced CosPlace, enhancing GeM by integrating it with a linear projection layer, achieving remarkable performance for VPR. Ali-beyĀ et al.Ā [4] introduced MixVPR, an all-MLP aggregation technique that achieved state-of-the-art performance on multiple benchmarks. Another facet of VPR is the training procedure, where the focus is on the loss functionĀ [3, 2, 12, 31].
@@
Transformer-based VPR. The Transformer architecture was initially introduced for natural language processingĀ [48], and later adapted to Vision Transformers (ViT) for computer vision applicationsĀ [20]. In VPR, they have recently been used as backbone combined with MixVPRĀ [27, 26] or NetVLADĀ [51] resulting in a performance boost compared to CNN backbones. Furthermore, AnyLocĀ [29] used features extracted from a foundation model, called DINOv2Ā [38], combined with unsupervised aggregation methods such as VLAD, resulting in notable performance on a wide range of benchmarks.
@@
Recently, two-stage retrieval strategy has gained prominence in Transformers-based VPR. It starts with global retrieval using global descriptors to identify top candidates, followed by a computationally-intensive reranking phase that refines the results based on local features. In this context, WangĀ et al.Ā [49] introduced TransVPR, which uses CNNs for feature extraction and Transformers for attention fusion to create global image descriptors, with additional patch-level descriptors for geometric verification. On the other hand, ZhangĀ et al.Ā [57] proposed ETR, a Transformer-based reranking technique that uses a CNN backbone for local and global descriptors. Their method leverages cross-attention between the local features of two images to cast the reranking as classification. Recently, ZhuĀ et al.Ā [62] proposed a unified framework integrating global retrieval and reranking, solely using transformers. For global retrieval, it employs the class token and utilizes other image tokens as local features for the reranking module. These two-stage techniques showed great performance, but at the expense of more computation and memory overheads. However, their global retrieval performance is still very limited compared to one-stage methods such as MixVPRĀ [4] and CosPlaceĀ [11]. In this work, we propose, Bag-of-Queries (BoQ), a Transformer-based aggregation technique for global retrieval, that outperforms existing state-of-the-art without relying on reranking. This makes BoQ particularly suitable for applications where computational resources are limited, yet high accuracy and efficiency are essential.
@@ 3 Methodology
In visual place recognition, effective aggregation of local features is crucial for accurate and fast global retrieval. To address this, we propose the Bag-of-Queries (BoQ) technique, a novel aggregation architecture that is end-to-end trainable and surprisingly simple, as depicted in Fig.Ā 2.
Our technique rely on Multi-Head Attention (MHA) mechanismĀ [48], which takes three inputs, queries (), keys () and values (), and linearly project them into multiple parallel heads. The output, for each head, is computed as follows:
@@
This ensures that the final descriptor combines information aggregated at different levels of the network, forming a rich and hierarchical representation. The final step involves reducing the dimensionality of the output using one or two successive linear projections with weights and , similar to the approach inĀ [3, 4].
This is followed by -normalization, to bring the global descriptor to the unit hypersphere, which optimizes the similarity search. The training is performed using pair-based (or triplet-based) loss functions that are widely used in the literatureĀ [3].
Relations to other methods. The Bag-of-Queries (BoQ) approach bears conceptual resemblance to Detection Transformer (DETR) modelĀ [14] in its use of object queries. However, BoQ is fundamentally different, as its global queries are exclusively used to extract and aggregate local feature information from the input; these queries do not contribute directly to the final representation, as demonstrated by the absence of any residual connection between the global queries and the output of the cross-attention. This is in contrast to DETR, where the object queries are integral part to the output, upon which object detection and classification is done.
In comparison to NetVLADĀ [6], which employs a set of learned cluster centers, and aggregates local features by calculating the weighted sum of their distances to each center. BoQ, leverages cross-attention between its learned queries and the input local feature to dynamically assess their relevance.
@@ 4.1 Datasets
Our experiments are performed on existing benchmarks that present a wide range of real-world challenges for VPR systems. Tab.Ā 1 provides a brief summary of these benchmarks. MapillarySLS (MSLS)Ā [53] was collected using dashcams and features a broad range of viewpoints and lighting changes, testing the systemās adaptability to typical variations in ground-level VPR. Pitts250kĀ [46] and Pitts30k datasets, extracted from Google Street View, exhibit significant changes in viewpoint, which tests the systemās ability to maintain recognition across varying angles and perspectives. AmsterTimeĀ [54] presents a unique challenge with historical grayscale images as queries and contemporary color images as references, covering temporal variations over several decades. EynshamĀ [18] consists entirely of grayscale images, adding complexity due to the absence of color information. NordlandĀ [44] dataset was collected over four seasons, using a camera mounted on the front of a train. It encompasses scenes ranging from snowy winters to sunny summers, presenting the challenge of extreme appearance changes due to seasonal variations. Note that two variants of Nordland are used in VPR benchmarking: one uses a subset of 2760 summer images as queries against the entire winter sequence as reference images (marked withĀ *), while the other uses the entire winter sequence as query images against the entire summer sequence as reference images (marked withĀ **). SVOXĀ [10] stands out with its comprehensive coverage of weather conditions, including overcast, rainy, snowy, and sunny images, to test adaptability to meteorological changes. Additionally, SVOX Night subset focuses on nocturnal scenes, a challenging scenario for most VPR systems.
@@
| MSLSĀ [53] | k | ||||
| Pitts250kĀ [46] | k | k | |||
| Pitts30kĀ [46] | k | k | |||
| AmsterTimeĀ [54] | |||||
| EynshamĀ [18] | k | k | |||
| Nordland*Ā [56] | k | ||||
| Nordland**Ā [44] | k | k | |||
| St-LuciaĀ [36] | |||||
| SVOXĀ [10] | k | k | |||
| SPEDĀ [56] | |||||
@@ 4.2 Implementation details
Architecture. For our experiments, we primarily employ pre-trained ResNet-50 backbones for feature extractionĀ [25]. Importantly, our proposed Bag-of-Queries (BoQ) technique is architecturally agnostic and seamlessly integrates with various CNNs or Vision Transformer (ViT) backbones. When employing a ResNet, we crop it at the second last residual to preserve a higher resolution of local features. Additionally, we freeze all but the last residual block and add a convolution to reduce the number of channels, which decreases the computational and memory footprint of the self-attention mechanism. Implementing the BoQ model is straightforward with popular deep learning frameworks such as PyTorchĀ [39] and TensorFlowĀ [1], both providing highly optimized implementations of self-attention and cross-attention, which are the basic blocks of our technique.
Training. We train BoQ following the standard framework of GSV-CitiesĀ [3], which provides a highly accurate dataset of k images depicting k places. For the loss function, we use Multi-Similarity lossĀ [50], as it has been shown to perform best for visual place recognitionĀ [3]. We use batches containing places, each depicted by images resulting in mini-batches of images. We use the AdamW optimizer with a learning rate of depending the batch-size, and a weight decay of . The learning rate is multiplied by after each epochs. Finally, we train for a maximum of epochs (including epochs of linear warmup) using images resized to . Note that while current techniques such as NetVLADĀ [6], CosPlaceĀ [61] and EigenPlacesĀ [12] train on images of size āeffectively triple the pixel count of our chosen resolutionāour BoQ model still achieves better performance, underscoring its effectiveness.
Evaluation metrics. We follow the same evaluation metric of existing literature [6, 30, 53, 56, 49, 11, 4, 3, 2, 12], where the recall@ is measured. Recall@ is defined as the percentage of query images for which at least one of the top- predicted reference images falls within a predefined threshold distance, which is typically meters. However, for the Nordland benchmark, which features aligned sequences, the threshold is frame counts for Nordland** and only frame for Nordland*.
@@ 4.3 Comparison with previous works
In this section, we present an extensive comparison of our proposed method, Bag-of-Queries (BoQ), with a wide range of existing state-of-the-art VPR aggregation techniques. This includes global average pooling (AVG)Ā [6], Generalized MeanĀ (GeM)Ā [41], NetVLADĀ [6] alongside its recent variants, SPE-NetVLADĀ [55] and GatedĀ NetVLADĀ [58]. Importantly, we compare against recent cutting-edge techniques, such as ConvAPĀ [3], CosPlaceĀ [61], EigenPlacesĀ [12] and MixVPRĀ [4], which currently hold the best scores in most existing benchmarks, setting a high standard for BoQ to measure up against in the field of VPR.
Moreover, despite BoQ being a global retrieval technique which does not employ reranking, we compare it with current state-of-the-art two-stage retrieval techniquesĀ [19, 13, 24, 49, 62]. These methods leverage geometric verification to enhance recall@k performance at the expense of increased computation and memory overhead.
Results analysis. In Tab.Ā 2, we present a comparative analysis across several challenging datasets, focusing on Recall@1 (R@1), Recall@5 (R@5), and Recall@10 (R@10). We also offer insight into the performance of each method with regard to the top-ranked retrieved images.
ā¢
@@
In Tab.Ā 3, we follow the same evaluation style proposed by BertonĀ et al.Ā [12], where benchmarks are categorized into multi-view datasets, where images are oriented in various angles, and frontal-view dataset where images are mostly forward facing. Top-1 retrieval performance (R@1) is reported for each technique.
ā¢
@@
ā¢
On SVOX, BoQ achieves best results under all conditions, especially on SVOX night, where it achieves R@1, outperforming the second best method (MixVPR) by percentage points. This highlights BoQās potential under low-light conditions.
Furthermore, we compare BoQ to two-stage retrieval techniques in Tab.Ā 4, by reporting performance on MSLS-val and Pitts30k-test. Although our technique does not perform reranking, its global retrieval performance surpasses that of existing two-stage techniques, including Patch-NetVLAD (2021), TransVPR (2022) and R2Former (2023). Importantly, this makes our approach more efficient in terms of memory and computation: over faster retrieval time compared to R2Former.